OpenStudy (anonymous):

4x-4 bigger than or equal to 2 x +10 bigger than or equal to 4

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

rewrite ur question again......it doesn't make any sense

OpenStudy (anonymous):

So...\[4x-4\ge2x+10\ge4\]? I assume solve for x (as best as one can, anyway)? \[2x-2\ge x+5\ge2\]\[2x-x-2\ge x-x+5\ge2-x\]\[x-2\ge5\ge2-x\]\[x-2 \ge 5 \ge -(x-2)\]\[|5| \le x-2\]\[5+2 \le x\]\[x \ge 7\]

OpenStudy (help!!!!):

you are correcto